OpenStudy (anonymous):

find the derivative of x+5/x+2, instead of -3/x+2^2, i got 3/x+2^2

myininaya (myininaya):

\[(\frac{x+5}{x+2})'=\frac{(x+5)'(x+2)-(x+2)'(x+5)}{(x+2)^2}\]

myininaya (myininaya):

\[=\frac{(1+0)(x+2)-(1+0)(x+5)}{(x+2)^2}\] \[=\frac{x+2-(x+5)}{(x+2)^2}\]

myininaya (myininaya):

\[=\frac{-3}{(x+2)^2}\]

OpenStudy (anonymous):

thanks! quick question. in any other problem like this, do you always sub in 1?

myininaya (myininaya):

derivative of x is 1 i didn't sub in 1 thats just what the derivative of x is